# Achieving Near-Optimal Regret for Bandit Algorithms with Uniform Last-Iterate Guarantee

### v1

@@ D.1 ULI guarantee for SE-MAB algorithm
Consider the successive elimination (SE-MAB) algorithm (e.g., Algorithm 3 in (Even-Dar et al., 2006)) in the MAB setting where ${\mathcal{A}}=[K]$. We present SE-MAB algorithm in Algorithm 6 and define some notations. Let $N_{a}(t)$ be the number of times that arm $a$ has been pulled before round $t$ and let $\widehat{\mu}_{a}(s)$ be the empirical mean of arm $a$ after $s\in\mathbb{N}$ number of plays. We define the confidence width for each $s\in\mathbb{N}$ as
@@
Following the standard trick (e.g., (Audibert et al., 2010)), we rewrite the empirical mean for every $a,n$ as $\widehat{\mu}_{a}(n)=\frac{1}{n}\sum_{i=1}^{n}R_{i,a}$ where $R_{i,a}$ is the reward of $i$-th pull of arm $a$. Let us fix both $s\in\mathbb{N}$ and $a\in[K]$. By Hoeffding’s inequality, with probability at least $1-\delta^{\prime}$
@@ D.2 ULI guarantee for PE-MAB
In MAB setting with ${\mathcal{A}}=[K]$, we further consider phased elimination algorithm (e.g., Exercise 6.8 of (Lattimore and Szepesvári, 2020)) shown in Algorithm 7 (called PE-MAB). The algorithm proceeds with phases $\ell=1,2,\ldots$, and each phase $\ell$ includes consecutive rounds, with an exponential increase. In phase $\ell$, the algorithm sequentially pulls every arm $a\in[K]$ for $m_{\ell}$ times where
@@ D.3 ULI guarantee for PE-linear
We consider a phased elimination algorithm (e.g., algorithm in Chapter 22 of (Lattimore and Szepesvári, 2020)) for linear bandits setting with a finite arm set ${\mathcal{A}}=[K]$. The algorithm proceeds with phases $\ell=1,2,\ldots$, and in each phase $\ell$, the algorithm first computes a design $\pi_{\ell}\in\Delta({\mathcal{A}}_{\ell})$ over all active arms where $\Delta({\mathcal{A}}_{\ell})$ is the set of all Radon probability measures over set ${\mathcal{A}}_{\ell}$. Rather than computing an exact design in (Lattimore and Szepesvári, 2020), we follow Lattimore et al. (2020) to compute a nearly-optimal design (19), which can be efficiently implemented. Then, PE-linear plays each arm $a\in{\mathcal{A}}_{\ell}$ for $m_{\ell}(a)$ times and updates the active arm set by using the estimates in this phase.
@@
By Eq. (20.2) of (Lattimore and Szepesvári, 2020), we have with probability at least $1-\delta/(2\ell^{2}K)$,
@@
The choice of $U_{\delta}(x)$ in Eq. (22) is slightly different from that of (Jamieson et al., 2014) because we consider for any $\delta\in(0,1)$ and they constrain the choice of $\delta$ in a more restricted range. The choice of $U_{\delta}(x)$ is motivated from another lemma of law of iterated logarithm, given in (Dann et al., 2017, Lemma F.1) with $\sigma^{2}=1/4$ as we here consider $[0,1]$-bounded rewards. Note that the concentration bounds in (Dann et al., 2017) apply for the conditionally subgaussian random variables, and one can get the same result for i.i.d. subgaussian random variables, by simply replacing the Doob’s maximal inequality by Hoeffding’s maximal inequality (see (Jamieson et al., 2014, Lemma 3)).
@@ E.2 Proof of Supporting Lemmas
Appendix F Omitted Details of Section 3.3
Our main objective in this section is to prove the following results.
Theorem F.1 (Restatement of Theorem 3.5).
@@
[…]
@@
[…]
We decompose the proof of Theorem 3.5 into two parts. In the first part, we first show that any algorithm with Condition. 3.4 for some $\delta\in(0,1)$ can achieve the ULI guarantee. Then, we only need to show that Exp3.P and GeometricHedge.P satisfy Condition. 3.4, which completes the proof of Theorem 3.5.
The following result suggests that showing an algorithm enjoys the ULI guarantee is reduced to show that this algorithm meets Condition. 3.4.
@@
For any fixed $\delta\in(0,1)$, if Algorithm 2 uses a base-algorithm that satisfies Condition. 3.4, then, it achieves the ULI guarantee that
@@
With Theorem F.2 in hand, our next objective is to show that Exp3.P and GeometricHedge.P are able to satisfy Condition. 3.4. For GeometricHedge.P with John’s exploration (which improves the original bound (Bartlett et al., 2008)), Lee et al. (2021) have already shown that it achieves the condition, and the rest result will show that Exp3.P also holds it. The full proof can be found in Appendix F.2.
@@
In MAB setting, for any $\delta\in(0,1)$, Exp3.P with given arm set ${\mathcal{A}}$, satisfies Condition. 3.4 with $g(\delta)=\mathcal{O}\left(|{\mathcal{A}}|\log\left(|{\mathcal{A}}|/\delta\right)\right)$.
The key that Exp3.P as well as GeometricHedge.P can fulfill Condition. 3.4 is because it adds a small amount of probability for uniform exploration to each arm. Therefore, the importance-weighted (IW) estimator can be lower-bounded and the term $\left|\sum_{t}(\widehat{\ell}_{t,a}-\ell_{t,a})\right|$ will not be too large.
Apart from Exp3.P for MAB and GeometricHedge.P for linear bandits, Lee et al. (2021) show that refined version of SCRiBLe (Lee et al., 2020) can be used as a base-algorithm for Algorithm 2 to achieve the ULI guarantee for linear bandits, but the ULI guarantee is inferior to that of GeometricHedge.P (i.e., inferior $F_{\text{ULI}}$ w.r.t. $d$).
F.1 Proof of Theorem F.2
@@
If Algorithm 2 accepts a base-algorithm which satisfies Condition. 3.4 as an input, then, with probability at least $1-\frac{3}{5}\delta$, for all $m\in\mathbb{N}$ and $a\in{\mathcal{A}}_{m}$,
@@
We first consider a fixed phase $m$. Since the base-algorithm satisfies Condition. 3.4, if the base-algorithm runs for consecutive $T_{m}$ and active arm set ${\mathcal{A}}_{m}\subseteq{\mathcal{A}}$, then, with probability at least $1-\frac{3}{5}\delta^{\prime}$, for all $a\in{\mathcal{A}}_{m}$
@@
Recall from the second bullet of Condition. 3.4 that
@@
Since for any fixed $\delta\in(0,1)$, $g(\delta/x^{2})$ is monotonically increasing for $x\geq 1$ (recall Condition. 3.4) and Lemma F.12 gives $m(t)\leq\log_{2}(t+1)$, we have
@@
According to the first bullet of Condition. 3.4, the proof is complete. ∎
F.2 Proof of Proposition F.3
In this section, we prove that Exp3.P (Auer et al., 2002b) meets Condition. 3.4. In our setting, arm set is ${\mathcal{A}}=[K]$, and the loss $\ell_{t,a}$ generated by each arm $a$ at each round $t$ is assumed to be $[0,1]$-bounded.
@@
Lemma F.14 (Exercise 5.15 of (Lattimore and Szepesvári, 2020)).
@@
As for GeometricHedge.P, we consider an improved version in (Lee et al., 2021) which uses John’s exploration and shows that $g(\delta)=d\log(K/\delta)$. Note that this is slightly different from $d\log(K\log_{2}T/\delta)$ presented in (Lee et al., 2021). The extra $\log_{2}T$ term is caused by Lemma 2 of (Bartlett et al., 2008), a concentration inequality for martingales, but one can invoke Lemma F.17 to avoid it. By applying such a function $g(\delta)$ into Theorem F.2, we obtain the claimed result.

### v2

@@ D.1 ULI guarantee for SE-MAB algorithm
Consider the successive elimination (SE-MAB) algorithm (e.g., Algorithm 3 in [Even-Dar et al., 2006]) in the MAB setting where ${\mathcal{A}}=[K]$. We present SE-MAB algorithm in Algorithm 6 and define some notations. Let $N_{a}(t)$ be the number of times that arm $a$ has been pulled before round $t$ and let $\widehat{\mu}_{a}(s)$ be the empirical mean of arm $a$ after $s\in\mathbb{N}$ number of plays. We define the confidence width for each $s\in\mathbb{N}$ as
@@
       Play an arm $A_{t}=\operatorname*{\arg\!\min}_{a\in{\mathcal{A}}_{t}}N_{a}(t)$ and observe reward $X_{t,A_{t}}$. Update counter $N_{a}(t+1)=N_{a}(t)+1$ for $A_{t}=a$ and $N_{a}(t+1)=N_{a}(t)$ for all $a\in{\mathcal{A}}_{t}-\{A_{t}\}$. Update empirical means $\widehat{\mu}_{A_{t}}(N_{A_{t}}(t+1))=\frac{\widehat{\mu}_{A_{t}}(N_{A_{t}}(t))\cdot N_{A_{t}}(t)+X_{t,A_{t}}}{N_{A_{t}}(t)+1}$ and $\widehat{\mu}_{a}(N_{a}(t+1))=\widehat{\mu}_{a}(N_{a}(t))$ for all $a\in{\mathcal{A}}_{t}-\{A_{t}\}$. Update confidence width $0pt(N_{a}(t+1))$ based on Eq. (11) for all $a\in{\mathcal{A}}_{t}$. Update bad arm set ${\mathcal{B}}_{t}$ as
${\mathcal{B}}_{t}=\left\{a\in{\mathcal{A}}_{t}:\exists j\in{\mathcal{A}}_{t}\text{ such that }\widehat{\mu}_{j}(N_{j}(t+1))-0pt(N_{j}(t+1))-\widehat{\mu}_{a}(N_{a}(t+1))-0pt(N_{a}(t+1))>0\right\}.$
Update active arm set ${\mathcal{A}}_{t+1}={\mathcal{A}}_{t}-{\mathcal{B}}_{t}$.
@@
Following the standard trick (e.g., [Audibert et al., 2010]), we rewrite the empirical mean for every $a,n$ as $\widehat{\mu}_{a}(n)=\frac{1}{n}\sum_{i=1}^{n}R_{i,a}$ where $R_{i,a}$ is the reward of $i$-th pull of arm $a$. Let us fix both $s\in\mathbb{N}$ and $a\in[K]$. By Hoeffding’s inequality, with probability at least $1-\delta^{\prime}$
@@ D.2 ULI guarantee for PE-MAB
In MAB setting with ${\mathcal{A}}=[K]$, we further consider phased elimination algorithm (e.g., Exercise 6.8 of [Lattimore and Szepesvári, 2020]) shown in Algorithm 7 (called PE-MAB). The algorithm proceeds with phases $\ell=1,2,\ldots$, and each phase $\ell$ includes consecutive rounds, with an exponential increase. In phase $\ell$, the algorithm sequentially pulls every arm $a\in[K]$ for $m_{\ell}$ times where
@@
       Play every arm $a\in{\mathcal{A}}_{\ell}$ for $m_{\ell}$ times and observe corresponding rewards. Update empirical means $\left\{\widehat{\mu}_{a}(m_{\ell})\right\}_{a\in{\mathcal{A}}_{\ell}}$ only using samples from phase $\ell$. Update active arm set ${\mathcal{A}}_{\ell+1}$ as
${\mathcal{A}}_{\ell+1}={\mathcal{A}}_{\ell}-\left\{a\in{\mathcal{A}}_{\ell}:\max_{j\in{\mathcal{A}}_{\ell}}\widehat{\mu}_{j}(m_{\ell})-\widehat{\mu}_{a}(m_{\ell})>2^{-\ell}\right\}.$
@@ D.3 ULI guarantee for PE-linear
       Find a design $\pi_{\ell}\in\Delta({\mathcal{A}}_{\ell})$ with
$\max_{a\in{\mathcal{A}}_{\ell}}\left\|{a}\right\|_{G_{\ell}^{-1}}^{2}\leq 2d,\quad\text{and}\quad\left|\text{supp}(\pi_{\ell})\right|\leq 4d\log\log(d)+16,$ (13)
where $G_{\ell}=\sum_{a\in{\mathcal{A}}_{\ell}}\pi_{\ell}(a)aa^{T}$. Play every arm $a\in{\mathcal{A}}_{\ell}$ for $m_{\ell}(a)=\left\lceil\pi_{\ell}(a)m_{\ell}\right\rceil$ times and observe corresponding rewards. Update the empirical estimate as
$\widehat{\theta}_{\ell}=V_{\ell}^{-1}\sum_{t\in{\mathcal{T}}_{\ell}}A_{t}X_{t,A_{t}},\quad\text{where}\quad V_{\ell}=\sum_{a\in{\mathcal{A}}_{\ell}}m_{\ell}(a)aa^{\top}.$ (14)
Update active arm set
${\mathcal{A}}_{\ell+1}={\mathcal{A}}_{\ell}-\left\{a\in{\mathcal{A}}_{\ell}:\max_{b\in{\mathcal{A}}_{\ell}}\left\langle{\widehat{\theta}_{\ell},b-a}\right\rangle>2^{-\ell+1}\right\}.$
We consider a phased elimination algorithm (e.g., algorithm in Chapter 22 of [Lattimore and Szepesvári, 2020]) for linear bandits setting with a finite arm set ${\mathcal{A}}=[K]$. The algorithm proceeds with phases $\ell=1,2,\ldots$, and in each phase $\ell$, the algorithm first computes a design $\pi_{\ell}\in\Delta({\mathcal{A}}_{\ell})$ over all active arms where $\Delta({\mathcal{A}}_{\ell})$ is the set of all Radon probability measures over set ${\mathcal{A}}_{\ell}$. Rather than computing an exact design in [Lattimore and Szepesvári, 2020], we follow Lattimore et al. [2020] to compute a nearly-optimal design (13), which can be efficiently implemented. Then, PE-linear plays each arm $a\in{\mathcal{A}}_{\ell}$ for $m_{\ell}(a)$ times and updates the active arm set by using the estimates in this phase.
@@
By Eq. (20.2) of [Lattimore and Szepesvári, 2020], we have with probability at least $1-\delta/(2\ell^{2}K)$,
@@
The choice of $U_{\delta}(x)$ in Eq. (16) is slightly different from that of [Jamieson et al., 2014] because we consider for any $\delta\in(0,1)$ and they constrain the choice of $\delta$ in a more restricted range. The choice of $U_{\delta}(x)$ is motivated from another lemma of law of iterated logarithm, given in [Dann et al., 2017, Lemma F.1] with $\sigma^{2}=1/4$ as we here consider $[0,1]$-bounded rewards. Note that the concentration bounds in [Dann et al., 2017] apply for the conditionally subgaussian random variables, and one can get the same result for i.i.d. subgaussian random variables, by simply replacing the Doob’s maximal inequality by Hoeffding’s maximal inequality (see [Jamieson et al., 2014, Lemma 3]).
@@
       Play an arm
$A_{t}=\operatorname*{\arg\!\max}_{a\in{\mathcal{A}}}\left\{\widehat{\mu}_{a}(N_{a}(t))+U_{\delta}\left(N_{a}(t)\right)\right\},$
where $U_{\delta}(N_{a}(t))$ is given in Eq. (16). Update counters $N_{a}(t+1)=N_{a}(t)+1$ for $A_{t}=a$ and $N_{a}(t+1)=N_{a}(t)$ for all $a\neq A_{t}$. Update empirical means $\widehat{\mu}_{A_{t}}(N_{A_{t}}(t+1))=\frac{\widehat{\mu}_{A_{t}}(N_{A_{t}}(t))\cdot N_{A_{t}}(t)+X_{t,A_{t}}}{N_{A_{t}}(t)+1}$ and $\widehat{\mu}_{a}(N_{a}(t+1))=\widehat{\mu}_{a}(N_{a}(t))$ for all $a\in{\mathcal{A}}-\{A_{t}\}$.
@@ E.2 Proof of Supporting Lemmas
E.3 Traditional Optimistic Algorithms Fail to Achieve ULI Guarantee
We will show that the traditional optimistic algorithms fail to achieve the ULI guarantee due to the $\log t$ term in the bonus. At a high level, $\log t$ is increasing with time and forces the algorithm to play a bad arm indefinitely when $t$ evolves. We provide the formal proof in the following:
We prove the claim by contradiction. Consider a $K$-armed bandit instance with at least one suboptimal arm, and let $\Delta>0$ be the minimum arm gap. Suppose there exists an optimism-based algorithm with $\log t$ in bonus term that can achieve the ULI guarantee in this setting. Then, for some fixed $\delta\in(0,1)$, with probability $\geq 1-\delta$, for all $t\in\mathbb{N}$, $\Delta_{t}\leq F_{ULI}(\delta,t)$. Based on Definition 2.5, we have $\lim_{t\to\infty}F_{ULI}(\delta,t)=0$ and $F_{ULI}(\delta,t)$ is monotonically decreasing w.r.t. $t$ after a threshold. Thus, $\exists t_{0}\in\mathbb{N}$ such that $F_{ULI}(\delta,t)<\Delta$ for all $t\geq t_{0}$. In other words, the algorithm cannot play any suboptimal arm after $t_{0}$-th round. Recall that the bonus term is $\sqrt{\log t/N_{a}(t)}$ where $N_{a}(t)$ is the number of plays of arm $a$ before round $t$. For any suboptimal arm $a$, $N_{a}(t)$ should not increase after $t_{0}$-th round, but $\log t$ keeps increasing. This leads the bonus of arm $a$ goes to infinity, which will incur a play of arm $a$ at a round after $t_{0}$. This makes a contradiction.
Appendix F Achieving ULI by Adversarial Bandit Algorithms
F.1 Meta-algorithm Enabling Adversarial Algorithms to Achieve ULI
Algorithm 10 Meta-algorithm towards ULI
Input: $\delta\in(0,1)$, finite arm set ${\mathcal{A}}$, and base-algorithm $\mathtt{Alg}$ with function $g(\delta)$ that satisfies Condition. F.1.
Initialize: ${\mathcal{A}}_{1}=[K]$.
for $m=1,2,\ldots$ do
      432 Set duration $T_{m}=\left\lceil 576g_{m}2^{m}\right\rceil$ where $g_{m}=g\left(\frac{\delta}{2m^{2}}\right)$. Run $\mathtt{Alg}$ with active arm set ${\mathcal{A}}_{m}$ for $T_{m}$ rounds and construct ^$\widehat{\ell}_{t,a}=\frac{\ell_{t,a}\cdot\mathbb{I}{\left\{{A_{t}=a}\right\}}}{p_{t,a}}$ for all $t,a$ in this phase. Update active arm set ${\mathcal{A}}_{m+1}={\mathcal{A}}_{m}-{\mathcal{B}}_{m}$ where:
${\mathcal{B}}_{m}=\left\{a\in{\mathcal{A}}_{m}:\sum_{s\in{\mathcal{T}}_{m}}\left(\widehat{\ell}_{s,a}-\widehat{\ell}_{s,k}\right)>7\sqrt{g_{m}T_{m}}\right\},$ (20)
and ${\mathcal{T}}_{m}$ is a set that contains all rounds in phase $m$ and $k$ is the empirical best arm:
$k\in\operatorname*{\arg\!\min}_{a\in{\mathcal{A}}_{m}}\sum_{s\in{\mathcal{T}}_{m}}\widehat{\ell}_{s,a}$
In this subsection, we propose a meta-algorithm shown in Algorithm 10 which enables any high-probability adversarial algorithm, with a mild condition, to achieve the ULI guarantee. Then, we show that existing high-probability adversarial bandit algorithms naturally meet this condition in both MAB and linear bandit settings. For notational simplicity, we follow the convention of adversarial analysis to use loss $\ell_{t,a}=1-X_{t,a}$. Note that all algorithms in this subsection require $[0,1]$-boundedness assumption on loss.
Condition F.1.
In fact, many high-probability adversarial bandit algorithms naturally meet Condition. F.1, but require stronger analysis. In particular, we show that Exp3.P [Auer et al., 2002b] for MAB holds this condition with $g(\delta)=\mathcal{O}\big{(}|{\mathcal{A}}|\log(|{\mathcal{A}}|/\delta)\big{)}$ (omitted proof can be found in Appendix F) and Lee et al. [2021] show that GeometricHedge.P [Bartlett et al., 2008] with John’s exploration meets the condition for linear bandits with $g(\delta)=\mathcal{O}\big{(}d\log(|{\mathcal{A}}|/\delta)\big{)}$. Hence, feeding those algorithms to Algorithm 10 yields the following:
Theorem F.2.
@@ F.1 Meta-algorithm Enabling Adversarial Algorithms to Achieve ULI
    $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{K\log\left(\delta^{-1}K\log(t+1)\right)}\big{)}.$
@@
    $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{d\log\left(\delta^{-1}K\log(t+1)\right)}\big{)}.$
Theorem F.2 suggests that Algorithm 10 enables Exp3.P and GeometricHedge.P to achieve near-optimal ULI guarantees for MAB and linear bandits, respectively. Moreover, their ULI guarantees are as good as those of conventional elimination-based algorithms (refer to Theorem 3.2).
Our main objective in this section is to prove the following results.
Theorem F.3 (Restatement of Theorem F.2).
For any fixed $\delta\in(0,1)$, if Algorithm 10 uses
    $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{K\log\left(\delta^{-1}K\log(t+1)\right)}\big{)}.$
    $F_{\text{ULI}}(\delta,t)=\mathcal{O}\big{(}t^{-\frac{1}{2}}\sqrt{d\log\left(\delta^{-1}K\log(t+1)\right)}\big{)}$
We decompose the proof of Theorem F.2 into two parts. In the first part, we first show that any algorithm with Condition. F.1 for some $\delta\in(0,1)$ can achieve the ULI guarantee. Then, we only need to show that Exp3.P and GeometricHedge.P satisfy Condition. F.1, which completes the proof of Theorem F.2.
The following result suggests that showing an algorithm enjoys the ULI guarantee is reduced to show that this algorithm meets Condition. F.1.
@@
For any fixed $\delta\in(0,1)$, if Algorithm 10 uses a base-algorithm that satisfies Condition. F.1, then, it achieves the ULI guarantee that
@@
With Theorem F.4 in hand, our next objective is to show that Exp3.P and GeometricHedge.P are able to satisfy Condition. F.1. For GeometricHedge.P with John’s exploration (which improves the original bound [Bartlett et al., 2008]), Lee et al. [2021] have already shown that it achieves the condition, and the rest result will show that Exp3.P also holds it. The full proof can be found in Appendix F.3.
@@
In MAB setting, for any $\delta\in(0,1)$, Exp3.P with given arm set ${\mathcal{A}}$, satisfies Condition. F.1 with $g(\delta)=\mathcal{O}\left(|{\mathcal{A}}|\log\left(|{\mathcal{A}}|/\delta\right)\right)$.
The key that Exp3.P as well as GeometricHedge.P can fulfill Condition. F.1 is because it adds a small amount of probability for uniform exploration to each arm. Therefore, the importance-weighted (IW) estimator can be lower-bounded and the term $\left|\sum_{t}(\widehat{\ell}_{t,a}-\ell_{t,a})\right|$ will not be too large.
Apart from Exp3.P for MAB and GeometricHedge.P for linear bandits, Lee et al. [2021] show that refined version of SCRiBLe [Lee et al., 2020] can be used as a base-algorithm for Algorithm 10 to achieve the ULI guarantee for linear bandits, but the ULI guarantee is inferior to that of GeometricHedge.P (i.e., inferior $F_{\text{ULI}}$ w.r.t. $d$).
F.2 Proof of Theorem F.4
@@
If Algorithm 10 accepts a base-algorithm which satisfies Condition. F.1 as an input, then, with probability at least $1-\frac{3}{5}\delta$, for all $m\in\mathbb{N}$ and $a\in{\mathcal{A}}_{m}$,
@@
We first consider a fixed phase $m$. Since the base-algorithm satisfies Condition. F.1, if the base-algorithm runs for consecutive $T_{m}$ and active arm set ${\mathcal{A}}_{m}\subseteq{\mathcal{A}}$, then, with probability at least $1-\frac{3}{5}\delta^{\prime}$, for all $a\in{\mathcal{A}}_{m}$
@@
Recall from the second bullet of Condition. F.1 that
@@
Since for any fixed $\delta\in(0,1)$, $g(\delta/x^{2})$ is monotonically increasing for $x\geq 1$ (recall Condition. F.1) and Lemma F.14 gives $m(t)\leq\log_{2}(t+1)$, we have
@@
According to the first bullet of Condition. F.1, the proof is complete. ∎
F.3 Proof of Proposition F.5
In this section, we prove that Exp3.P [Auer et al., 2002b] meets Condition. F.1. In our setting, arm set is ${\mathcal{A}}=[K]$, and the loss $\ell_{t,a}$ generated by each arm $a$ at each round $t$ is assumed to be $[0,1]$-bounded.
@@
       Play an arm $A_{t}\in[K]$ from distribution $p_{t}=[p_{t,1},\ldots,p_{t,K}]$ and observe loss $\ell_{t,A_{t}}$ where
$p_{t,a}=(1-\gamma)\frac{w_{t,a}}{W_{t}}+\frac{\gamma}{K}\quad\text{where}\quad W_{t}=\sum_{a\in[K]}w_{t,a}.$ (31)
Update $w_{t+1,a}=w_{t,a}\exp(-\eta\widetilde{\ell}_{t,a})$ for all $a\in[K]$ with
~$\widetilde{\ell}_{t,a}=\widehat{\ell}_{t,a}-\frac{2\beta}{p_{t,a}},\quad\text{where}\quad\widehat{\ell}_{t,a}=\frac{\ell_{t,a}B_{t,a}}{p_{t,a}},\text{ and }B_{t,a}=\mathbb{I}\left\{A_{t}=a\right\}.$ (32)
@@
Lemma F.16 (Exercise 5.15 of [Lattimore and Szepesvári, 2020]).
@@
As for GeometricHedge.P, we consider an improved version in [Lee et al., 2021] which uses John’s exploration and shows that $g(\delta)=d\log(K/\delta)$. Note that this is slightly different from $d\log(K\log_{2}T/\delta)$ presented in [Lee et al., 2021]. The extra $\log_{2}T$ term is caused by Lemma 2 of [Bartlett et al., 2008], a concentration inequality for martingales, but one can invoke Lemma F.19 to avoid it. By applying such a function $g(\delta)$ into Theorem F.4, we obtain the claimed result.